# Characteristic initial value problems for the Einstein-Maxwell-scalar field equations in spherical symmetry

### v2

@@ Abstract
The characteristic initial boundary problem is discussed in spherical symmetry for the Einstein-Maxwell-scalar field equations. It is formulated for an affine-null metric and the resulting field equations are cast into a hierarchical system of partial differential equations. The initial boundary value problem for a family of null hypersurfaces is specified for a timelike-null foliation at the central geodesic of spherical symmetry as well as for a double-null foliation where the corresponding boundary is a null hypersurface. For the latter, two distinct boundary value formulations arise – one where the null boundary has zero Misner-Sharp mass and another one where the corresponding Misner-Sharp mass is nonzero. As an application, the nonextremal and the extremal Reissner-Nordström solution in null coordinates for a charged black hole and the Fisher-Janis-Newman-Winicour solution are derived.
@@ I Introduction
The choice of spacetime foliation with correspondingly adapted coordinates to solve the Einstein field equations determines whether the resulting partial differential equations are either a complete set of hyperbolic partial differential equations or a mixed hyperbolic/elliptic system. The hypersurfaces of the hyperbolic part of this system along which the information travels are the characteristic surfaces. Characteristic surfaces in General Relativity may be for example null cones along which gravitational waves propagate or null hypersurfaces tied to the spacetime’s causal structure. Solution methods for Einstein equations using the geometric information of the characteristic are commonly referred to as characteristic methods. The employment of characteristic methods to study the field equations of General Relativity very often brought a deeper understanding to various aspects of the theory. Three important examples are the discovery of the Bondi mass loss formula by Bondi, Sachs and collaborators Bondi (1960); Bondi et al. (1962); Sachs (1962a), the first long term stable numerical evolutions of a black hole Gómez et al. (1998) via computer simulations and recently the proof of incorrectness of the third law of black hole thermodynamics Kehle and Unger (2022). The Bondi mass loss formula was found in an asymptotic patch of a spacetime where the radial coordinate parametrizes the null rays of outgoing null cones. The long term stable black hole simulations were possible by numerically representing the spacetime manifold with a countable number of finite size patches discretized by finite difference operators adapted to either in- or outgoing null hypersurfaces. The technical base of the disproof of the third law was the novel technique called “characteristic gluing” Aretakis et al. (2021), where different characteristic (null) hypersurfaces are stitched together using gauge invariant quantities. These three examples have in common that the patch(es) is (are) charted with at least one null coordinate, i.e. a coordinate constant along a null hypersurface, and that boundary information at one side of the patch needs to be provided to archive the physical result. For the Bondi mass loss case, this boundary data is the specification of an asymptotic inertial observer, an initial value for the mass and the so-called news function which is the retarded time derivative of the gravitational shear. For the black hole simulation, the boundary data were the specification of black hole data at an inner world tube excising the black hole singularity, where the first patch of the numerically discretized null hypersurface is located. While in the disproof of the third law, the boundary data between the patches are specification of quantities in Minkowski space and the horizon of a Schwarzschild/Reissner-Nordström black hole. The difference for the latter with respect to the former two is that the employed initial boundary value problem is based on the two intersecting null hypersurfaces instead of only one null hypersurface. A further difference is that in the last example the Einstein equations are solved with a coupling to a charged scalar field, while the first two concerned the full Einstein equations in vacuum. Although in spherical symmetry, the coupling to the scalar field captures dynamical features and adds the necessary degrees of freedom so that the discussion of a characteristic initial boundary value problem (CIBVP) can be made without loosing too much structural information.
@@
Here, we investigate characteristic methods for Einstein equations minimally coupled to a charged, massless and necessary complex scalar field in spherical symmetry. In contrast to the numerous works Angelopoulos et al. (2024); Kehle and Unger (2024a, b, 2022); Kehle and Van de Moortel (2024) related to characteristic gluing Aretakis et al. (2021) as well as the studies of formation of sub-extremal black hole formation of Reisner-Nordström black holes Murata et al. (2013); Gelles and Pretorius (2025) with characteristic methods using a double null foliation, we will investigate the initial boundary value problem in affine null metric coordinates Winicour (2013). The affine null metric formulation is build on a foliation with respect to a family of null hypersurfaces and has the advantage that a simple evolution algorithm can be found, which is beneficial for numerical work. For this evolution scheme, the field equations can be cast into one set of hierarchical hypersurface equations and one first order transport equation with respect to the chosen null coordinate. The set of derived equations is similar to a previously used one for a real scalar field Crespo et al. (2019) having the additional advantage that it can be used to investigate black hole spacetimes beyond the formation of an event horizon and track the apparent horizon and the scalar field inside the black hole Mädler et al. (2024). This interior black hole evolution is possible because the hypersurface-evolution equations remain regular where the expansion rate of the outgoing null rays vanishes, i.e. at an apparent horizon. Here, we extend and modify the scheme of Crespo et al. (2019) such that this regularisation is maintained if the scalar field is complex and a Maxwell field is added to the energy momentum tensor. We will also show that for the derived set of equations, CIBVPs can be formulated either on the central world line in spherical symmetry or by two different versions of double null CIBVPs at a null boundary. As an example for these cases, we discuss the asymptotic CIBVP showing that the entire asymptotic behaviour of a charged scalar field is governed by the specification of an initial value for the scalar field monopole, mass, charge and a free function - the scalar field news function. We further derive the subextremal and extremal Reisner-Nordström black hole as well as the Fisher-Janis-Newman-Winicour (FJNW) solution in the affine null-metric coordinate chart. Those solution turn out to be expressible by simple rational functions.
@@
In Sec. II, the basic field equations and additional geometric quantities for the Einstein-Maxwell scalar field equations in spherical symmetry and affine null metric coordinates are presented. The regularised set of hypersurface-evolution equations following from the field equations being the CIBVPs is given in Sec. III. The CIBVP in an asymptotic region is presented in III.1, the one at the central geodesic is formulated in III.2, while the two different CIBVPs with a null boundary surface are given in Sec. III.3. In Sec. IV, we discuss the properties for asymptotic timelike Killing vectors, which are needed for the derivation of the Reisner-Nordström black hole in Sec. V.1 and the FJNW solution in Sec. V.2. Sec. VI contains a summary and conclusion of the work.
@@
We use geometric units with $c=1$ and the Einsteinian gravitational constant is $\kappa=8\pi G$. The metric signature is $+2$, capital latin indices $A,B,C,...$ are two dimensional with values $x^{A}=(x^{2},x^{3})$ and lowercase latin indices are four dimensional taking values $x^{a}=(w,\lambda,x^{A})$. The covariant derivative of the metric is $\nabla_{a}$ and partial derivatives are denoted with the comma notation $\frac{\partial f}{\partial x^{a}}=f_{,a}$. Calligraphic capital letters, $\mathcal{J},\mathcal{L},\mathcal{C},...$ are always assumed complex fields, while Roman capital letters, like $Q,Z,V,L,\ldots$ are purely real. Numeral subscripts in parenthesis, like $V_{(1)}$, correspond to expansion coefficients of a power series evaluated at $\lambda=0$ with the number in the parenthesis referring to the power $\lambda^{k}$. Numeral subscripts in a bracket, like $\Phi_{[1]}$, correspond to expansion coefficients of an asymptotic power series and the number in the bracket refers to the order of the inverse power $1/\lambda^{k}$.
@@ II Field equations for the Einstein-Maxwell-scalar field equations
In a four dimensional spacetime, consider the Einstein-Hilbert action $S$ given by
$S=\int d^{4}x\sqrt{-g}\left[\frac{g^{ab}R_{ab}}{2\kappa}-\frac{F_{ab}F^{ab}}{16\pi}-\frac{1}{2}(\overline{\mathcal{D}^{a}\Phi})(\mathcal{D}_{a}\Phi)\right]$ (1)
where $\kappa=8\pi G$ is the Einsteinian gravitational constant, $\sqrt{-g}$ the determinant of the metric $g_{ab}$ with well defined inverse $g^{ab}$ and corresponding Levi-Civita covariant derivative $\nabla_{c}$, $R_{ab}$ is the Ricci tensor wrt $g_{ab}$, $F_{ab}=\nabla_{a}A_{b}-\nabla_{b}A_{a}$ is a Maxwell field determined by the 4-covector $A_{b}$, $\mathcal{D}_{c}=\nabla_{c}+iqA_{c}$ a gauge invariant derivative for the massless complex scalar field $\Phi$ that is coupled to the Maxwell field with the gauge charge $q$. The overbar denotes complex conjugation. From the variation of the action with respect to $g^{ab}$, $A_{a}$, $\bar{\Phi}$ and $\Phi$, we can derive the trace reversed Einstein equations, $E_{ab}$, the Maxwell equations, $E_{a}$, and the scalar wave equations $E$ and $\bar{E}$, which are respectively
@@
$-\kappa\left[\frac{1}{2}(\overline{\mathcal{D}_{a}\Phi})(\mathcal{D}_{b}\Phi)+\frac{1}{2}(\overline{\mathcal{D}_{b}\Phi})(\mathcal{D}_{a}\Phi)\right]$
$E_{a}:=$ 00 0 $=\nabla_{a}F^{a}_{\;b}+4\pi j_{b}$ (3)
$E:=$ 00 0 $={\mathcal{D}^{a}}\mathcal{D}_{a}\Phi$ (4)
$\bar{E}:=$ 00 0 $=\overline{\mathcal{D}^{a}\mathcal{D}_{a}\Phi}$ (5)
@@
$j_{b}=\frac{iq}{2}\left(\overline{\Phi}\mathcal{D}_{a}\Phi-\Phi\overline{\mathcal{D}_{a}\Phi}\right)$ (6)
@@
with $g_{AB}$ being the 2-metric of the spacelike hypersurface of the co-dimension 2 boundary $\Sigma=\partial\mathtt{V}$ where the integral curves of null vectors $\ell^{a}\partial_{a}$ and $n^{a}\partial_{a}$ orthogonal to $\Sigma$ intersect. Let us consider the metric
$ds^{2}=-V(w,\lambda)dw^{2}+2\epsilon dwd\lambda+r^{2}(w,\lambda)q_{AB}dx^{A}dx^{B}$ (8)
where $q_{AB}$ is any metric of a unit sphere [^2] In standard spherical coordinates $x^{A}=(\theta,\varphi)$, $q_{AB}=\mathrm{diag}(1,\sin^{2}\theta)$ with the three associated Killing vectors $\xi_{1}=\partial_{\phi}$, $\xi_{2}=\cos\phi\partial_{\theta}-\cot\theta\sin\phi\partial_{\phi}$ and $\xi_{3}=\sin\phi\partial_{\theta}+\cot\theta\cos\phi\partial_{\phi}$. In principle, this two metric could also be a flat metric, like ${q_{AB}=\mathrm{diag}(1,1)}$, or a unit hypersphere ${q_{AB}=\mathrm{diag}(1,\sinh^{2}\theta)}$. Of the field equations, only (17) would change by the substitution of the term +1 by either $+0$ or $-1$, respectively. The former may be interesting for plane wave front propagation, like pp waves, while the latter for hyperbolical black holes or fluids, see e.g. Herrera et al. (2020, 2021).. An understanding of the the binary parameter $\epsilon$ can be found in flat space where $V=1$ and $r=\lambda$, then for $\epsilon=-1$ the coordinate surfaces $w=const$ represents outgoing null hypersurfaces $u$, while for $\epsilon=1$ they are ingoing null hypersurfaces $v=const$.
@@
where we used $\epsilon^{2}=1$. Furthermore, the form of the metric (8) remains invariant under the coordinate changes
@@
For the metric we define two null vectors
@@
whose associated one forms are $\ell_{a}dx^{a}=\epsilon dw$ and $n_{a}dx^{a}=\epsilon(V/2)dw-d\lambda$ and which are orthogonal to one another $n^{a}\ell_{a}=-1$. The null vector field $\ell^{a}\partial_{a}$ is tangent to affine parameterized null geodesics meaning $\ell^{a}\nabla_{a}\ell^{a}=0$, while the vector field $n^{a}\partial_{a}$ is in general tangent to non-affine parametrized null geodesics where
@@
The non-zero components of $E_{ab}$ are
$(E_{ww})$ ::: $0=r_{,ww}-\frac{V(r^{2}V_{,\lambda})_{,\lambda}}{4r}-\frac{\epsilon}{2}\left(r_{,w}V_{,\lambda}-r_{,\lambda}V_{,w}\right)\;,+\frac{1}{2}\kappa r\left[|\Phi_{,w}|^{2}+\frac{VQ^{2}}{8\pi r^{4}}+q^{2}\alpha^{2}|\Phi|^{2}+iq\alpha\Bigl{(}\Phi\bar{\Phi}_{,w}-\bar{\Phi}\Phi_{,w}\Bigr{)}\right]\;,$
$(E_{w\lambda})\;$ ::: $0=r_{,w\lambda}+\frac{\epsilon}{4r}(r^{2}V_{,\lambda})_{,\lambda}-\frac{1}{2}\kappa r\left[\frac{\epsilon Q^{2}}{8\pi r^{4}}-\frac{1}{2}[\overline{\Phi}_{,w}\Phi_{,\lambda}+\overline{\Phi}_{,\lambda}\Phi_{,w}]-\frac{iq\alpha}{2}(\Phi\overline{\Phi}_{,\lambda}-\overline{\Phi}\Phi_{,\lambda})\right]\;,$
@@
$(q^{AB}E_{AB})\;$ ::: $0=-\left(Vrr_{,\lambda}+2\epsilon rr_{,w}\right)_{,\lambda}+1-\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}}$ (17)
and those of $E_{a}$
@@
In spherical symmetry the Misner-Sharp mass Misner and Sharp (1964)
$M=\frac{r}{2}(1-g^{ab}r_{,a}r_{b})=\frac{r}{2}\Bigl{(}1-2\epsilon r_{,w}r_{,\lambda}-Vr_{,\lambda}^{2}\Bigr{)}$ (21)
is an invariant concept of quasilocal mass. Assuming asymptotically flat spacetimes, its asymptotic limit
$m_{B}:=\lim_{\lambda\rightarrow\infty}M$ (22)

### v3

@@ Abstract
The characteristic initial boundary problem is discussed in spherical symmetry for the Einstein-Maxwell-scalar field equations. It is formulated for an affine-null metric and the resulting field equations are cast into a hierarchical system of partial differential equations. The initial boundary value problem for a family of null hypersurfaces is specified for a timelike-null foliation at the central geodesic of spherical symmetry as well as for a double-null foliation where the corresponding boundary is a null hypersurface. For the latter, two distinct boundary value formulations arise; one where the null boundary has zero Misner-Sharp mass and another one where the corresponding Misner-Sharp mass is nonzero. As an application, the nonextremal and the extremal Reissner-Nordström solution in null coordinates for a charged black hole and the Fisher-Janis-Newman-Winicour solution are derived.
@@ I Introduction
The choice of spacetime foliation with correspondingly adapted coordinates to solve the Einstein field equations determines whether the resulting partial differential equations are either a complete set of hyperbolic partial differential equations or a mixed hyperbolic/elliptic system. The hypersurfaces of the hyperbolic part of this system along which the information travels are the characteristic surfaces. Characteristic surfaces in general relativity may be for example null cones along which gravitational waves propagate or null hypersurfaces tied to the spacetime’s causal structure. Solution methods for Einstein equations using the geometric information of the characteristic are commonly referred to as characteristic methods. The employment of characteristic methods to study the field equations of general relativity very often brought a deeper understanding to various aspects of the theory. Three important examples are the discovery of the Bondi mass loss formula by Bondi, Sachs and collaborators Bondi (1960); Bondi et al. (1962); Sachs (1962a), the first long term stable numerical evolutions of a black hole Gómez et al. (1998) via computer simulations and recently the proof of incorrectness of the third law of black hole thermodynamics Kehle and Unger (2022). The Bondi mass loss formula was found in an asymptotic patch of a spacetime where the radial coordinate parametrizes the null rays of outgoing null cones. The long term stable black hole simulations were possible by numerically representing the spacetime manifold with a countable number of finite size patches discretised by finite difference operators adapted to either in- or outgoing null hypersurfaces. The technical base of the disproof of the third law was the novel technique called “characteristic gluing” Aretakis et al. (2021), where different characteristic (null) hypersurfaces are stitched together using gauge invariant quantities. These three examples have in common that the patch(es) is (are) equipped with at least one null coordinate, i.e. a coordinate constant along a null hypersurface, and that boundary information at one side of the patch needs to be provided to achieve the physical result. For the Bondi mass loss case, this boundary data are the specification of an asymptotic inertial observer, an initial value for the mass and the so-called news function which is the retarded time derivative of the gravitational shear. For the black hole simulation, the boundary data were the specifications of black hole data at an inner world tube excising the black hole singularity, where the first patch of the numerically discretised null hypersurface is located. While in the disproof of the third law, the boundary data between the patches are specifications of quantities in Minkowski space and the horizon of a Schwarzschild/Reissner-Nordström black hole. The difference for the latter with respect to the former two is that the employed initial boundary value problem is based on the two intersecting null hypersurfaces instead of only one null hypersurface. A further difference is that in the last example the Einstein equations are solved with a coupling to a charged scalar field, while the first two concerned the full Einstein equations in vacuum. In spherical symmetry, the coupling to the scalar field captures dynamical features and adds the necessary degrees of freedom so that the discussion of a characteristic initial boundary value problem (CIBVP) can be made without losing too much structural information.
@@
Here, we investigate characteristic methods for Einstein equations minimally coupled to a charged, massless and necessary complex scalar field in spherical symmetry. In contrast to the numerous works Angelopoulos et al. (2024); Kehle and Unger (2024a, b, 2022); Kehle and Van de Moortel (2024) related to characteristic gluing Aretakis et al. (2021) as well as the studies of formation of subextremal Reissner-Nordström black holes Murata et al. (2013); Gelles and Pretorius (2025) with characteristic methods using a double-null foliation, we will investigate the CIBVP in affine-null metric coordinates Winicour (2013). The affine-null metric formulation is built on a foliation with respect to a family of null hypersurfaces and has the advantage that a simple evolution algorithm can be found, which is beneficial for numerical work. As we will demonstrate, for this evolution scheme, the field equations can be cast into one set of hierarchical hypersurface equations and one first order transport equation with respect to the chosen null coordinate. The set of derived equations is similar to a previously used one for a real scalar field Crespo et al. (2019) having the additional advantage that it can be used to investigate black hole spacetimes beyond the formation of an event horizon and track the apparent horizon and the scalar field inside the black hole Mädler et al. (2024). This interior black hole evolution is possible because the hypersurface-evolution equations remain regular where the expansion rate of the outgoing null rays vanishes, i.e. at an apparent horizon.
In fact, with an extension of Crespo et al. (2019), it is possible to investigate the Choptuik phenomenon Choptuik (1993) for charged scalar fields by globally foliating spacetime to include null infinity on the numerical grid through coordinate compactification111 It was shown in Mädler and Gallo (2025) that the conformal Einstein-scalar field equations in affine-null coordinates Friedrich (1986); Penrose (1963) are equivalent to corresponding physical Einstein-scalar field equations, provided the latter employ a compactified affine parameter and suitable normalizations of the metric and scalar field. Thus, null infinity can be represented on a numerical grid using a compactified coordinate.. Traditionally, such investigations have been carried out only within the 3+1 formulation of the Einstein-Maxwell-scalar field equations Gundlach and Martin-Garcia (1996), which does not usually allow this. These studies can also address the formation process of charged, spherically symmetric black holes. Indeed, the CIVBP presented below can be used to numerically test the mathematical results of Kehle (2019), particularly regarding the disproof of the third law of black hole thermodynamics, that is, as a consequence of Kehle and Unger (2022), extremal black holes should form in finite time. In a numerical simulation, this would imply identifying an extremal black hole in the output of a simulation where the black hole forms within finite simulation time. In this context, it is significantly simpler to study the formation of extremal black holes in the Einstein-Maxwell-scalar field system in spherical symmetry rather than to study the formation of extremal Kerr black holes. This is because, in the former, the simulations can be performed with a single spherically symmetric dynamical field (i.e., the scalar field) , whereas the latter requires at least axial symmetry for the Einstein equations, involving two dynamical fields (i.e. the gravitational wave degrees of freedom).
Affine-null coordinates can also be useful for constructing outgoing null coordinate systems that not only describe the exterior of black holes, but are also sufficiently regular to describe the region inside the event horizon. Coordinate systems of this kind for the Schwarzschild case include, for example, the Israel coordinates Israel (1966), which were later generalized to include a broader family of metrics inGallo et al. (2021). Other null coordinate systems have been constructed numerically for the Kerr black hole in Argañaraz and Moreschi (2021), and used to describe the evolution of a scalar field penetrating the black hole interior in Argañaraz and Moreschi (2022). A fully analytical construction of an affine-null coordinate system to describe slowly rotating Kerr black holes is also possible, as demonstrated in Mädler and Gallo (2023). The general formalism presented in this work for spherically symmetric spacetimes accommodates such regular affine-null coordinate systems as a particular case, thereby allowing for the study of matter fields radiating in the exterior region while simultaneously tracking their evolution as they penetrate into the interior of black holes.
In this work, we extend and modify the scheme of Crespo et al. (2019) so that the regularization of the fields involved in the hierarchical affine-null system of equations is preserved when the scalar field is complex and a Maxwell field is included in the energy-momentum tensor. We will also show that for the derived set of equations, CIBVPs can be formulated either on the central world line in spherical symmetry or by two different versions of double-null CIBVPs at a null boundary. As an example for these cases, we discuss the asymptotic CIBVP showing that the entire asymptotic behavior of a charged scalar field is governed by the specification of an initial value for the scalar field monopole, mass, charge and a free function - the scalar field news function. We further derive the subextremal and extremal Reissner-Nordström black hole as well as the Fisher-Janis-Newman-Winicour (FJNW) solution in the affine-null-metric coordinate chart. Those solutions turn out to be expressible by simple rational functions.
@@
In Sec. II, the basic field equations and additional geometric quantities for the Einstein-Maxwell-scalar field equations in spherical symmetry and affine-null metric coordinates are presented. The regularized set of hypersurface-evolution equations following from the field equations being the CIBVPs is given in Sec. III. The CIBVP in an asymptotic region is presented in Sec. III.1, the one at the central geodesic is formulated in III.2, while the two different CIBVPs with a null boundary surface are given in Sec. III.3. In Sec. IV, we discuss the properties for asymptotic timelike Killing vectors, which are needed for the derivation of the Reissner-Nordström black hole in Sec. V.1 and the FJNW solution in Sec. V.2. Section VI contains a summary and conclusion of the work.
@@
We use geometric units with $c=1$ and the Einstein gravitational constant is $\kappa=8\pi G$. The metric signature is $+2$, capital Latin indices $A,B,C,\ldots$ are two dimensional with values $x^{A}=(x^{2},x^{3})$ and lowercase Latin indices are four dimensional taking values $x^{a}=(w,\lambda,x^{A})$. The covariant derivative of the metric is $\nabla_{a}$ and partial derivatives are denoted with the comma notation $\frac{\partial f}{\partial x^{a}}=f_{,a}$. Calligraphic capital letters, $\mathcal{J},\mathcal{L},\mathcal{C},...$ are always assumed complex fields, while Roman capital letters, such as $Q,Z,V,L,\ldots$ are purely real. Numeral subscripts in parentheses, like $V_{(1)}$, correspond to expansion coefficients of a power series evaluated at $\lambda=0$ with the number in the parentheses referring to the power $\lambda^{k}$. Numeral subscripts in brackets, such as $\Phi_{[1]}$, correspond to expansion coefficients of an asymptotic power series and the number in the bracket refers to the order of the inverse power $1/\lambda^{k}$.
@@ II Field equations for the Einstein-Maxwell-scalar field equations
In a four-dimensional spacetime, consider the Einstein-Hilbert action $S$ given by
$S=\int d^{4}x\sqrt{-g}\left[\frac{g^{ab}R_{ab}}{2\kappa}-\frac{F_{ab}F^{ab}}{16\pi}-\frac{1}{2}(\overline{\mathcal{D}^{a}\Phi})(\mathcal{D}_{a}\Phi)\right],$ (1)
where $\kappa=8\pi G$ is the Einsteinian gravitational constant, $\sqrt{-g}$ is the determinant of the metric $g_{ab}$ with a well defined inverse $g^{ab}$ and a corresponding Levi-Civita covariant derivative $\nabla_{c}$, $R_{ab}$ is the Ricci tensor with respect to $g_{ab}$, $F_{ab}=\nabla_{a}A_{b}-\nabla_{b}A_{a}$ is a Maxwell field determined by the four-covector $A_{b}$, and $\mathcal{D}_{c}=\nabla_{c}+iqA_{c}$ is a gauge invariant derivative for the massless complex scalar field $\Phi$ that is coupled to the Maxwell field with the gauge charge $q$. The overbar denotes complex conjugation. From the variation of the action with respect to $g^{ab}$, $A_{a}$, $\bar{\Phi}$ and $\Phi$, we can derive the trace reversed Einstein equations, $E_{ab}$, the Maxwell equations, $E_{a}$, and the scalar wave equations $E$ and $\bar{E}$, which are, respectively
@@
$-\kappa\left[\frac{1}{2}(\overline{\mathcal{D}_{a}\Phi})(\mathcal{D}_{b}\Phi)+\frac{1}{2}(\overline{\mathcal{D}_{b}\Phi})(\mathcal{D}_{a}\Phi)\right],$
$E_{a}:=$ 00 0 $=\nabla_{a}F^{a}_{\;b}+4\pi j_{b},$ (3)
$E:=$ 00 0 $={\mathcal{D}^{a}}\mathcal{D}_{a}\Phi,$ (4)
$\bar{E}:=$ 00 0 $=\overline{\mathcal{D}^{a}\mathcal{D}_{a}\Phi},$ (5)
@@
$j_{b}=\frac{iq}{2}\left(\overline{\Phi}\mathcal{D}_{a}\Phi-\Phi\overline{\mathcal{D}_{a}\Phi}\right),$ (6)
@@
with $g_{AB}$ being the two-metric of the spacelike hypersurface of the co-dimension two boundary $\Sigma=\partial\mathtt{V}$, which is the intersection of the integral curves of null vectors $\ell^{a}\partial_{a}$ and $n^{a}\partial_{a}$ orthogonal to $\Sigma$. Consider the metric
$ds^{2}=-V(w,\lambda)dw^{2}+2\epsilon dwd\lambda+r^{2}(w,\lambda)q_{AB}dx^{A}dx^{B},$ (8)
where $q_{AB}$ is any metric of a unit sphere [^3] In standard spherical coordinates $x^{A}=(\theta,\varphi)$, ${q_{AB}=\mathrm{diag}(1,\sin^{2}\theta)}$ with the three associated Killing vectors $\xi_{1}=\partial_{\phi}$, $\xi_{2}=\cos\phi\partial_{\theta}-\cot\theta\sin\phi\partial_{\phi}$ and $\xi_{3}=\sin\phi\partial_{\theta}+\cot\theta\cos\phi\partial_{\phi}$. In principle, this two metric could also be a flat metric, like ${q_{AB}=\mathrm{diag}(1,1)}$, or a unit hypersphere ${q_{AB}=\mathrm{diag}(1,\sinh^{2}\theta)}$. Of the field equations, only (17) would change by the substitution of the term +1 by either $+0$ or $-1$, respectively. The former may be interesting for plane wave front propagation, such as pp waves, while the latter for hyperbolical black holes or fluids (see e.g. Herrera et al. (2020, 2021)).. An understanding of the binary parameter $\epsilon$ can be found in flat space, where $V=1$ and $r=\lambda$, then for $\epsilon=-1$ the coordinate surfaces $w=$const represent outgoing null hypersurfaces $u$=const, while for $\epsilon=1$ they are ingoing null hypersurfaces $v=$const.
@@
where we used $\epsilon^{2}=1$. The form of the metric (8) remains invariant under the coordinate changes
@@
For the metric, we define two null vectors
@@
whose associated one forms are $\ell_{a}dx^{a}=\epsilon dw$ and $n_{a}dx^{a}=\epsilon(V/2)dw-d\lambda$ satisfying $n^{a}\ell_{a}=-1$. The null vector field $\ell^{a}\partial_{a}$ is tangent to affinely parametrized null geodesics meaning $\ell^{a}\nabla_{a}\ell^{a}=0$, while the vector field $n^{a}\partial_{a}$ is in general tangent to nonaffinely parametrized null geodesics, where
@@
The nonzero components of $E_{ab}$ are
$(E_{ww})$ ::: $0=r_{,ww}-\frac{V(r^{2}V_{,\lambda})_{,\lambda}}{4r}-\frac{\epsilon}{2}\left(r_{,w}V_{,\lambda}-r_{,\lambda}V_{,w}\right)+\frac{1}{2}\kappa r\left[|\Phi_{,w}|^{2}+\frac{VQ^{2}}{8\pi r^{4}}+q^{2}\alpha^{2}|\Phi|^{2}+iq\alpha\Bigl{(}\Phi\bar{\Phi}_{,w}-\bar{\Phi}\Phi_{,w}\Bigr{)}\right]\;,$
$(E_{w\lambda})\;$ ::: $0=r_{,w\lambda}+\frac{\epsilon}{4r}(r^{2}V_{,\lambda})_{,\lambda}-\frac{1}{2}\kappa r\left[\frac{\epsilon Q^{2}}{8\pi r^{4}}-\frac{1}{2}\left(\overline{\Phi}_{,w}\Phi_{,\lambda}+\overline{\Phi}_{,\lambda}\Phi_{,w}\right)-\frac{iq\alpha}{2}(\Phi\overline{\Phi}_{,\lambda}-\overline{\Phi}\Phi_{,\lambda})\right]$ (15)
@@
$(q^{AB}E_{AB})\;$ ::: $0=-\left(Vrr_{,\lambda}+2\epsilon rr_{,w}-\lambda\right)_{,\lambda}-\frac{\kappa}{8\pi}\frac{Q^{2}}{r^{2}},$ (17)
and those of $E_{a}$ are
@@
In spherical symmetry the Misner-Sharp mass Misner and Sharp (1964),
$M=\frac{r}{2}(1-g^{ab}r_{,a}r_{b})=\frac{r}{2}\Bigl{(}1-2\epsilon r_{,w}r_{,\lambda}-Vr_{,\lambda}^{2}\Bigr{)},$ (21)
is an invariant concept of quasilocal mass. Assuming asymptotically flat spacetimes, its asymptotic limit,
$m_{B}:=\lim_{\lambda\rightarrow\infty}M,$ (22)